OpenStudy (anonymous):

write a polynomial function in factored form with the given zeros -5,-2,0

OpenStudy (mertsj):

Could you solve this equation: (x+6)(x-4)=0

OpenStudy (anonymous):

you change (x+6) to x-6=0 so x=-6 right; then x-4=0 so x=4. Or am I completely off?

OpenStudy (mertsj):

No. That is very good. Now what you need to notice is that the ANSWER is x = 4 and the FACTOR is x-4.

OpenStudy (mertsj):

The other ANSWER is x = -6 and the FACTOR is x+6. Are you with me?

OpenStudy (anonymous):

yes I got ya so far

OpenStudy (mertsj):

So, in your problem, it says the answers are -5, -2 and 0. So what would the factors be?

OpenStudy (anonymous):

(x+5)(x+2)(x-0)

OpenStudy (anonymous):

(x+5)(x+2)(x-0)

OpenStudy (mertsj):

Put f(x) = on that and you have it!!!